# Majorization for probability distributions, column stochastic matrices and their linear preservers

### v1

@@ Abstract
In this paper we investigate majorization for probability distributions and column stochastic matrices. We show that majorizations in general can be reduced to these sets. We characterize linear operators that preserve majorization for probability distributions, and show their equivalence to operators preserving vector majorization. Our main result provides a complete characterization of linear preservers of strong majorization for column stochastic matrices, revealing a richer structure of preservers, than in the standard setting. As a prerequisite to this characterization, we solve the problem of characterizing linear preservers of majorization for zero-sum vectors, which yields a new structural insight into the classical results of Ando and of Li and Poon.
@@ 1 Introduction and main results
We write $A\geq O$ (resp. $v\geq 0$) when all entries of the matrix $A$ (resp. the vector $v$) are nonnegative.
Vectors in $\mathbb{R}^{n}$ are treated as column vectors and identified with corresponding $n$-tuples. We denote:
@@
  • For $v\in\mathbb{R}^{n}$ and $\mathcal{I}\subseteq\mathcal{N}$, $v_{\mathcal{I}}$ denotes the vector comprised of entries of $v$ indexed by $\mathcal{I}$
  • $v^{+}$ denotes the sum of all positive entries of $v\in\mathbb{R}^{n}$
@@
  • $\mathbbold{1}^{n}=\{v\in\mathbb{R}^{n}:e^{t}v=1\text{ and }v\geq 0\}$, the set of probability distributions.
@@
Majorization provides a powerful mathematical framework for comparing degrees of disorder in vectors and matrices, with applications ranging from linear algebra and economic inequality measures to quantum information theory and statistical experiments.
For a vector $x\in{\mathbb{R}}^{n}$ we let $x^{\downarrow}$ denote the permutation of its entries in the non-increasing order. For $a,b\in{\mathbb{R}}^{n}$ we say that $a$ is majorized by $b$, denoted by $a\preceq b$, if
@@
A matrix is called row (resp. column) stochastic if it is nonnegative and each row (resp. column) sum is equal to $1$. A matrix is called doubly stochastic, if it is both row and column stochastic. The set of all $n\times n$ row, column or doubly stochastic matrices is denoted by $\Omega_{n}^{row},\Omega^{col}_{n},\Omega_{n}$, respectively. Similarly, the set of all $n\times m$ column stochastic matrices is denoted by $\Omega^{col}_{n,m}$.
@@
[17, Theorem I.2.A.2] The elements of the set $P(n)$ of permutation matrices are the extreme points of $\Omega_{n}$. Moreover, $\Omega_{n}$ is the convex hull of matrices in $P(n)$.
@@
In this paper we investigate majorization for column stochastic matrices. In particular, we study vector majorization for probability distributions (single-column stochastic matrices). These problems are motivated by several aspects. It turns out that majorization in general can be reduced to majorization for column stochastic matrices and probability distributions in the vector case. On the other hand, majorizations of these objects are very important in applications, for instance, in Quantum Information Theory and Theory of Statistical experiments.
Vector majorization is often used to compare vectors and probability distributions, offering insights into the concepts of order and disorder in various systems. The quantum mechanical analogue of a probability distribution is the density matrix. A density matrix $A$ is said to be more chaotic or less pure than the density matrix $B$, if the spectrum of $A$ is majorized by the spectrum of $B$. This notion is usually called Quantum Majorization or simply Matrix Majorization depending on the context. Spectra of density matrices are exactly probability distributions.
@@
Majorization for column stochastic matrices arises in the Theory of Statistical Experiments. In the discrete settings, statistical experiments are expressed as column stochastic matrices. The matrix majorization relation introduced in [5] turns out to be the criteria of one experiment being more informative than the other. Strong majorization, that we focus on, is a particular (stronger) case of this majorization. Further information can be found in [25, 4, 5, 6]. Note that in these references, compared to the present paper, mostly the transposed versions of the notions are used. In particular, the matrices are row-stochastic.
@@
A linear operator $\Phi$ on $\mathcal{M}_{n,m}$ preserves strong majorization (resp. preserves strong majorization on $\Omega^{col}_{n,m}$) if $A\preceq^{s}B$ implies $\Phi(A)\preceq^{s}\Phi(B)$ for any $A,B\in M_{n,m}$ (resp. for any $A,B\in\Omega^{col}_{n,m}$).
@@
The main result of the paper is Theorem 6.19 that characterizes linear operators preserving strong majorization for column stochastic matrices. Despite the fact that strong majorization in general can also be reduced to column stochastic matrices, and despite the simplicity of the reduction, we obtain a richer structure of the preservers.
Notably, this problem turns out to be more complicated than other similar matrix majorization preserver problems in the following sense. A matrix operator on $M_{n,m}$ can be decomposed in a standard way into $m^{2}$ linear operators on ${\mathbb{R}}^{n}$, see Definition 6.1. In the case of preservers of majorizations for real or for $(0,1)$-matrices, the $m^{2}$ components of the operator preserve the respective majorization in the vector case. This approach already gives a strong necessary condition which goes a long way toward solving the characterization problem.
In the case of column stochastic matrices, however, we cannot claim that the components of a preserver preserves majorization for $n\times 1$ column stochastic matrices, i.e. probability distributions. Indeed, as we see from Theorem 6.19 and Example 7.1, this is not the case. However, we find that these operators must preserve majorization for zero sum vectors. The concise characterization of such operators is found in Theorem 5.25. This result provides better insight into the nature of Ando’s classical operators, showing that they emerge as two particular cases of this natural and simple general structure.
The rest of the paper is organized as follows. Section 2 contains several additional notions of majorization and preliminary results. In Section 3 we show how majorization can be reduced to column stochastic matrices. Section 4 provides characterization of linear preservers of majorization for probability distributions. Linear preservers of strong majorization for column stochastic matrices are characterized in Section 6. Section 5 solves the prerequisite problem of characterizing linear preservers of majorization for zero sum vectors. Finally, Section 7 presents several concluding observations and open problems.
@@ 2 Preliminaries
In this paper we demonstrate that strong majorization can be reduced to the case of column stochastic matrices. Since the same approach works for directional majorization and to a lesser extent for weak majorization, we also investigate these related majorizations.
@@
  • For $a,b\in{\mathbb{R}}^{n}$ it is said $a\sim b$ if $a\preceq b\preceq a$;
  • For $A,B\in M_{n,m}$ it is said $A\sim^{s}B$ if $A\preceq^{s}B\preceq^{s}A$.
@@ 3 Reduction to column stochastic matrices
In this section we establish that when dealing with strong or directional majorization, we can always reduce the problem to column stochastic matrices. The same reduction applies to vector majorization. When dealing with weak majorization, we can make one of the matrices column stochastic (whichever we choose). Additionally, we specifically examine the role of $(0,1)$-matrices.
@@ 3.1 Strong majorization
Let $A\preceq^{s}B$. Then $A=QB$ for some $Q\in\Omega_{n}$. Observe that $B+ev^{t}=QA+Qev^{t}$. Therefore $A+ev^{t}\preceq^{s}B+ev^{t}$.
@@
Let $A,B\in M_{n,m}$ and $\lambda\in{\mathbb{R}}$. Then $A\preceq^{s}B$ if and only if $A+\lambda J\preceq B+\lambda J$.
We now show how the general case of matrix majorization can be reduced to column stochastic matrices.
@@
    Choose $\lambda\geq 0$ large enough so that the matrices $A_{1}=A+\lambda J$ and $B_{1}+\lambda J$ are nonnegative. By Corollary 3.3, $A\preceq^{s}B$ if and only if $A_{1}\preceq^{s}B_{1}$.
@@
    Define $v\in{\mathbb{R}}^{n}$ by $v^{t}=\frac{1}{n}(e^{t}-e^{t}B_{2})$ and let $A^{\prime}=A_{2}+ev^{t}$ and $B^{\prime}=B_{2}+ev^{t}$. Note that $v$ is nonnegative by the definition of $B_{2}$. Then $A^{\prime}$ and $B^{\prime}$ are also nonnegative.
@@
Clearly, Corollary 3.5 is also satisfied for Reduction 3.7. This reduction offers a more direct approach via diagonal scaling than Reduction 3.4. On the other hand, Reduction 3.4 is more flexible and provides a more convenient general expression for $e^{t}A^{\prime}$.
@@
Then $D=diag(14,8,20,2)$ and we obtain $A^{\prime}=\left(\begin{matrix}\frac{6}{14}&\frac{5}{8}&\frac{11}{20}&\frac{1}{2}\\ \\ \frac{8}{14}&\frac{3}{8}&\frac{9}{20}&\frac{1}{2}\end{matrix}\right)$ and $B^{\prime}=\left(\begin{matrix}\frac{10}{14}&\frac{1}{8}&\frac{7}{20}&\frac{1}{2}\\ \\ \frac{4}{14}&\frac{7}{8}&\frac{13}{20}&\frac{1}{2}\end{matrix}\right)$.
@@ 3.2 Weak majorization
However, if in Reduction 3.4, Item 3 we substitute $v=\frac{1}{n}(e^{t}-e^{t}B_{2})$ with $w=\frac{1}{n}(e^{t}-e^{t}A_{2})$, then $A^{\prime}$ becomes column stochastic, while $B^{\prime}$, in general, does not. Analogously, in Reduction 3.7 we can substitute $D=diag(e^{t}B_{1})$ with $diag(e^{t}A_{1})$.
@@ 3.3 Directional majorization
Let $A\preceq^{d}B$. Consider arbitrary $x\in{\mathbb{R}}^{m}$. Then there exists $Q\in\Omega_{n}$ such that $Ax=QBx$. Observe that $Bx+ev^{t}x=QAx+Qev^{t}x$. Therefore $(A+ev^{t})x\preceq(B+ev^{t})x$. As $x$ was arbitrary, we obtain $A+ev^{t}\preceq^{d}B+ev^{t}$.
@@ 3.5 $(0,1)$-matrices
Observe that $e^{t}A^{\prime}=e^{t}B^{\prime}$ and matrices $A^{\prime},B^{\prime}$ have no zero columns. Let $D=\mbox{{d}iag$\,$}(e^{t}A^{\prime})=\mbox{{d}iag$\,$}(e^{t}B^{\prime})\in M_{m}$. Then the matrix $D$ is invertible. Moreover, $\Theta(A^{\prime})=A^{\prime}D^{-1}$ and $\Theta(B^{\prime})=B^{\prime}D^{-1}$. Then, by Corollary 2.4, we obtain that $A^{\prime}\preceq^{d}B^{\prime}$ if and only if $\Theta(A)=\Theta(A^{\prime})\preceq^{d}\Theta(B^{\prime})=\Theta(B)$, while $A^{\prime}\preceq^{s}B^{\prime}$ if and only if $\Theta(A^{\prime})\preceq^{s}\Theta(B^{\prime})$. ∎
@@
Linear operators preserving strong majorization for $(0,1)$-matrices were characterized in [12]. Note, however, that the operator $\Theta$ is not linear. Indeed, a map with column-stochastic image can not be linear, since the zero matrix is not column stochastic.
@@ 4 Linear operators preserving majorization for probability distributions
In this section we characterize linear operators preserving majorization for probability distributions. As we show in the lemma below, such operators necessarily preserve majorization for $(0,1)$-vectors.
@@
The characterization of linear operators preserving majorization for $(0,1)$-matrices was obtained in [12]. We provide here only the following two results concerning the vector case. They will be necessary for Theorem 4.5
@@
It remains to consider the case $n=3$. Let $a=\begin{pmatrix}2\\ 1\\ 0\end{pmatrix},b=\begin{pmatrix}2\\ 0\\ 1\end{pmatrix}$. Then $\frac{1}{3}a\sim\frac{1}{3}b$ and these vectors are probability distributions. Therefore $\phi$ preserves the majorization relation $\frac{1}{3}a\sim\frac{1}{3}b$. As a consequence, $\phi(a)\sim\phi(b)$. Then $\phi$ preserves vector majorization by Corollary 4.4. ∎
5 Linear operators preserving majorization for the zero sum vectors
In this section we characterize the following types of linear operators. Recall that $\mathbbold{0}^{n}$ is the set of all real zero-sum vectors. This characterization turns out to be crucial for linear operators preserving majorization for column stochastic matrices.

### v2

# Majorizations for probability distributions, column stochastic matrices and their linear preservers

@@ Abstract
In this paper, we study majorization for probability distributions and column stochastic matrices. We show that majorizations in general can be reduced to the aforementioned sets. We characterize linear operators that preserve majorization for probability distributions, and show their equivalence to operators preserving vector majorization. Our main result provides a complete characterization of linear preservers of strong majorization for column stochastic matrices, revealing a richer structure of preservers than in the standard setting. As a prerequisite to this characterization, we solve the problem of characterizing linear preservers of majorization for zero-sum vectors, which yields a new structural insight into the classical results of Ando and of Li and Poon.
@@ 1 Introduction and main results
We write $A\geq 0$ (resp. $v\geq 0$) if every entry of the matrix $A$ (resp. the vector $v$) is nonnegative.
Vectors in $\mathbb{R}^{n}$ are treated as column vectors and identified with the corresponding $n$-tuples. We denote:
@@
  • For $v\in\mathbb{R}^{n}$ and $\mathcal{I}\subseteq\mathcal{N}$, $v_{\mathcal{I}}$ denotes the vector comprising the entries of $v$ indexed by $\mathcal{I}$
  • $v^{+}$ denotes the sum of all positive entries of $v\in\mathbb{R}^{n}$.
@@
  • $\mathbbold{1}^{n}=\{v\in\mathbb{R}^{n}:e^{t}v=1\text{ and }v\geq 0\}$, the set of probability distributions, see Definition 1.4.
@@
Majorization is a powerful mathematical framework for comparing degrees of disorder in vectors and matrices, with applications ranging from linear algebra and economic inequality measures to quantum information theory and statistical experiments.
The main notion of this theory is vector majorization. For a vector $x\in{\mathbb{R}}^{n}$ we let $x^{\downarrow}$ denote the permutation of its entries in the non-increasing order. For $a,b\in{\mathbb{R}}^{n}$ we say that $a$ is majorized by $b$, denoted by $a\preceq b$, if
@@
A matrix is row stochastic (resp. column stochastic) if it is nonnegative and each row (resp. column) sum equals to $1$. A matrix is doubly stochastic if it is both row and column stochastic. The sets of all $n\times n$ row, column and doubly stochastic matrices are denoted by $\Omega_{n}^{row},\Omega^{col}_{n}$ and $\Omega_{n}$, respectively. Similarly, the set of all $n\times m$ column stochastic matrices is denoted by $\Omega^{col}_{n,m}$.
@@
[17, Theorem I.2.A.2] The elements of the set of permutation matrices $P(n)$ are the extreme points of $\Omega_{n}$. Moreover, $\Omega_{n}$ is the convex hull of matrices in $P(n)$.
@@
In this paper, we investigate majorization for column stochastic matrices. In particular, we study vector majorization for probability distributions.
Definition 1.4.
A vector $v\in{\mathbb{R}}^{n}$ is a probability distribution when $v\geq 0$ and $e^{t}v=1$. That is, a probability distribution is an $n\times 1$ column stochastic matrix.
Our investigations are motivated by several aspects. It turns out that majorization in general can be reduced to majorization for column stochastic matrices and probability distributions in the vector case. On the other hand, majorizations of these objects are very important in applications, for instance, in Quantum Information Theory and Theory of Statistical Experiments.
Vector majorization is often used to compare vectors and probability distributions, offering insights into the concepts of order and disorder in various systems. The quantum mechanical analogue of a probability distribution is the density matrix. A density matrix $A$ is said to be more chaotic or less pure than the density matrix $B$ if the spectrum of $A$ is majorized by the spectrum of $B$. This notion is usually called Quantum Majorization or simply Matrix Majorization, depending on the context. Spectra of density matrices are exactly probability distributions.
@@
Majorization for column stochastic matrices arises in the Theory of Statistical Experiments. In the discrete settings, statistical experiments are expressed as column stochastic matrices. The matrix majorization relation introduced in [5] turns out to be the criterion of one experiment being more informative than the other. Strong majorization that we focus on is a particular (stronger) case of this majorization. Further information can be found in [25, 4, 5, 6]. Note that in these references, compared to the present paper, mostly the transposed versions of the notions are used. In particular, the matrices are row stochastic.
@@
A linear operator $\Phi$ on $M_{n,m}$ preserves strong majorization (resp. preserves strong majorization on $\Omega^{col}_{n,m}$) if $A\preceq^{s}B$ implies $\Phi(A)\preceq^{s}\Phi(B)$ for any $A,B\in M_{n,m}$ (resp. for any $A,B\in\Omega^{col}_{n,m}$).
@@
The main result of the paper is Theorem 6.19 that characterizes linear operators preserving strong majorization for column stochastic matrices. Despite the fact that strong majorization in general can also be reduced to column stochastic matrices, and despite the simplicity of the reduction, we obtain a richer structure of preservers.
Notably, this problem turns out to be more complicated than other similar matrix majorization preserver problems in the following sense. A matrix operator on $M_{n,m}$ can be decomposed in a standard way into $m^{2}$ linear operators on ${\mathbb{R}}^{n}$, see Definition 6.1. In the case of preservers of majorizations for real or for $(0,1)$-matrices, the $m^{2}$ components of the operator leave invariant the respective majorization in the vector case. This approach already gives a strong necessary condition which goes a long way toward solving the characterization problem.
In the case of column stochastic matrices, however, we cannot claim that the components of a preserver leave invariant majorization for $n\times 1$ column stochastic matrices, i.e. probability distributions. Indeed, as we see from Theorem 6.19 and Example 6.21, this is not the case. However, we find that these operators must preserve majorization for zero-sum vectors. The concise characterization of such operators is found in Theorem 5.25. This result provides better insight into the nature of Ando’s classical operators, showing that they emerge as two particular cases of this natural and simple general structure.
The rest of the paper is organized as follows. Section 2 contains several additional notions of majorization and preliminary results. In Section 3 we show how majorization can be reduced to column stochastic matrices. Section 4 provides a characterization of linear preservers of majorization for probability distributions. Section 5 solves the prerequisite problem of characterizing linear preservers of majorization for zero-sum vectors. The characterization of linear preservers of strong majorization for column stochastic matrices is the main result of the paper that can be found in Section 6.
@@ 2 Preliminaries
In this paper, we demonstrate that strong majorization can be reduced to the case of column stochastic matrices. Since the same approach works for directional majorization and, to a slightly lesser extent, for weak majorization, we also investigate these related notions.
@@
  • For $a,b\in{\mathbb{R}}^{n}$ we write $a\sim b$ if $a\preceq b\preceq a$;
  • For $A,B\in M_{n,m}$ we write $A\sim^{s}B$ if $A\preceq^{s}B\preceq^{s}A$.
@@ 3 Reduction to column stochastic matrices
In this section, we establish that when dealing with strong or directional majorization, we can always reduce the problem to column stochastic matrices. The same reduction applies to vector majorization. When dealing with weak majorization, we can make either matrix in the majorization relation column stochastic. Additionally, we specifically examine the role of $(0,1)$-matrices.
@@ 3.1 Strong majorization
Let $A\preceq^{s}B$. Then $A=QB$ for some $Q\in\Omega_{n}$. Observe that $QB+Qev^{t}=A+ev^{t}$. Therefore $A+ev^{t}\preceq^{s}B+ev^{t}$.
@@
Let $A,B\in M_{n,m}$ and $\lambda\in{\mathbb{R}}$. Then $A\preceq^{s}B$ if and only if $A+\lambda J\preceq^{s}B+\lambda J$.
We now show how the general case of strong majorization can be reduced to column stochastic matrices.
@@
    Choose $\lambda\geq 0$ large enough so that the matrices $A_{1}=A+\lambda J$ and $B_{1}=B+\lambda J$ are nonnegative. By Corollary 3.3, $A\preceq^{s}B$ if and only if $A_{1}\preceq^{s}B_{1}$.
@@
    Define $v\in{\mathbb{R}}^{m}$ by $v^{t}=\frac{1}{n}(e^{t}-e^{t}B_{2})$ and let $A^{\prime}=A_{2}+ev^{t}$ and $B^{\prime}=B_{2}+ev^{t}$. Note that $v$ is nonnegative by the definition of $B_{2}$. Then $A^{\prime}$ and $B^{\prime}$ are also nonnegative.
@@
Clearly, Corollary 3.5 is also satisfied for Reduction 3.7. This reduction offers a more direct approach via the diagonal scaling than Reduction 3.4. On the other hand, Reduction 3.4 is more flexible and provides a more convenient general expression for $e^{t}A^{\prime}$.
@@
Then $D=\mbox{{d}iag$\,$}(14,8,20,2)$ and we obtain $A^{\prime}=\left(\begin{matrix}\frac{6}{14}&\frac{5}{8}&\frac{11}{20}&\frac{1}{2}\\ \\ \frac{8}{14}&\frac{3}{8}&\frac{9}{20}&\frac{1}{2}\end{matrix}\right)$ and $B^{\prime}=\left(\begin{matrix}\frac{10}{14}&\frac{1}{8}&\frac{7}{20}&\frac{1}{2}\\ \\ \frac{4}{14}&\frac{7}{8}&\frac{13}{20}&\frac{1}{2}\end{matrix}\right)$.
@@ 3.2 Weak majorization
However, if in Reduction 3.4, Item 3, we substitute $v=\frac{1}{n}(e^{t}-e^{t}B_{2})$ with $w=\frac{1}{n}(e^{t}-e^{t}A_{2})$, then $A^{\prime}$ becomes column stochastic, while $B^{\prime}$, in general, does not. Analogously, in Reduction 3.7 we can substitute $D=\mbox{{d}iag$\,$}(e^{t}B_{1})$ with diag ($\mbox{{d}iag$\,$}(e^{t}A_{1})$.
@@ 3.3 Directional majorization
Let $A\preceq^{d}B$. Consider arbitrary $x\in{\mathbb{R}}^{m}$. Then there exists $Q\in\Omega_{n}$ such that $Ax=QBx$. Observe that $QBx+Qev^{t}x=Ax+ev^{t}x$. Therefore $(A+ev^{t})x\preceq(B+ev^{t})x$. As $x$ was arbitrary, we obtain $A+ev^{t}\preceq^{d}B+ev^{t}$.
@@ 3.5 $(0,1)$-matrices
Observe that $e^{t}A^{\prime}=e^{t}B^{\prime}$ and matrices $A^{\prime},B^{\prime}$ do not have zero columns. Let $D=\mbox{{d}iag$\,$}(e^{t}A^{\prime})=\mbox{{d}iag$\,$}(e^{t}B^{\prime})\in M_{m}$. Then the matrix $D$ is invertible. Moreover, $\Theta(A^{\prime})=A^{\prime}D^{-1}$ and $\Theta(B^{\prime})=B^{\prime}D^{-1}$. Then, by Corollary 2.4, we obtain that $A^{\prime}\preceq^{d}B^{\prime}$ if and only if $\Theta(A)=\Theta(A^{\prime})\preceq^{d}\Theta(B^{\prime})=\Theta(B)$, while $A^{\prime}\preceq^{s}B^{\prime}$ if and only if $\Theta(A)\preceq^{s}\Theta(B)$. ∎
@@
Linear operators preserving strong majorization for $(0,1)$-matrices were characterized in [12]. Note, however, that the operator $\Theta$ is not linear. Indeed, a map with column stochastic image can not be linear, since the zero matrix is not column stochastic.
@@ 4 Linear operators preserving majorization for probability distributions
In this section, we characterize linear operators preserving majorization for probability distributions. As we show in the lemma below, such operators necessarily preserve majorization for $(0,1)$-vectors.
@@
The characterization of linear operators preserving majorization for $(0,1)$-matrices was obtained in [12]. We only provide here several relevant results that are concerned with the vector case.
@@
It remains to consider the case $n=3$. Let $a=\begin{pmatrix}2\\ 1\\ 0\end{pmatrix},b=\begin{pmatrix}2\\ 0\\ 1\end{pmatrix}$. Then $\frac{1}{3}a\sim\frac{1}{3}b$ and these vectors are probability distributions. Therefore $\phi$ preserves the equivalence $\frac{1}{3}a\sim\frac{1}{3}b$. As a consequence, $\phi(a)\sim\phi(b)$. Then $\phi$ preserves vector majorization by Corollary 4.4. ∎
5 Linear operators preserving majorization for the zero-sum vectors
In this section, we characterize linear preservers of majorization on $\mathbbold{0}^{n}$, see Definition 5.1. Recall that $\mathbbold{0}^{n}$ denotes the set of all real zero-sum vectors. This characterization turns out to be crucial for describing linear operators preserving majorization for column stochastic matrices.